a block of mass 2 kg is lying on the frictionless table .Aforce of 8N is applied on it for 12 s.Calculate its kinetic energy.

Grade:9

1 Answers

Vikas TU
9 years ago
From first law of eqn.
v = u + at
v = 0 + 8/2 x 12 = 48 m/s
 
  1. E = 0.5 x 2 x 48 x 48 J
  2. Now U can calculate..
Plz  approve!!!
 
for more :
facebook.com/pages/Ask-questions-PhysicsMathsChemistry/120940711383826